# IndicSuperTokenizer: An Optimized Tokenizer for Indic Multilingual LLMs

### v1

@@ Abstract
Tokenizers play a crucial role in determining the performance, training efficiency, and the inference cost of Large Language Models (LLMs). Designing effective tokenizers for multilingual LLMs is particularly challenging due to diverse scripts and rich morphological variation. While subword methods such as Byte Pair Encoding (BPE) are widely adopted, their effectiveness in multilingual settings remains underexplored. We present IndicSuperTokenizer, a tokenizer for Indic multilingual LLMs, that combines both subword and multi-word tokenization, along with language-specific pre-tokenization, leading to more linguistically aligned tokens and achieving a new state-of-the-art in fertility score. Evaluated across English, $22$ Indian languages and code data, our tokenizer improves the average fertility score by $39.5\%$ over LLaMA4 and by $18\%$ over Sutra (the current best). This translates to $44\%$ improvement in inference throughput over LLaMA4 while maintaining comparable performance on English and Indic benchmarks. We also present detailed ablations across tokenizer training data size, vocabulary size, merging techniques, and pre-tokenization strategies, demonstrating the robustness of our design choices.
@@ 1 Introduction
Large Language Models (LLMs) (touvron2023llama; grattafiori2024llama; abdin2025phi; guo2025deepseek; yang2025qwen3; team2025gemma) rely on the crucial step of tokenization, the process of converting raw text into discrete units called tokens. A key metric for evaluating tokenizers is the “fertility score” (or token-to-word ratio) (ali2024tokenizer) where, a lower fertility score is desirable due to more efficient (and hence cheaper) LLM training and inference. Among the many proposed approaches, subword tokenization schemes such as BPE (sennrich-etal-2016-neural), Unigram (kudo-2018-subword), WordPiece (song2021fast), and their byte-level extensions have become the de facto choice. However, tokenization remains an understudied topic within the LLM literature (dagan2024getting; mielke2021between), especially in multilingual settings (petrov2023language), where, skewed fertility scores across languages, often lead to concerns around fairness, high inference latency, cost and context size. With 22 constitutionally recognized languages [^1] https://en.wikipedia.org/wiki/Languages_with_official_recognition_in_India, these issues are especially pronounced for Indic languages comprising multiple scripts and a rich morphology. Our analysis suggests that tokenizers of popular multilingual tokenizers, largely designed for English, could produce fertility scores as high as 10.5 (LLaMA-4 tokenizer for Oriya; Table 22) for Indic languages, far worse than the near-ideal scores achieved for English. This leads to longer token sequences, higher compute overheads, and poor alignment with linguistic units like morphemes and compounds.
@@
Designing an efficient tokenizer involves making careful choices around the size of the vocabulary (of tokens), tokenizer training data, the tokenization approach, and, doing this across languages is nontrivial. In this work, we present IndicSuperTokenizer, an efficient tokenizer for Indic LLMs, that achieves state-of-the-art fertility scores across 22 Indic languages, English, and code. Our design choices are grounded in detailed ablations and our tokenizer combines linguistically grounded pre-tokenization with a two-stage subword–superword learning process (inspired from SuperBPE (liu2025superbpespacetravellanguage)), yielding a more compact and semantically faithful vocabulary. Figure 1 illustrates some examples where our approach avoids fragmenting common words or idiomatic phrases into unnatural subunits across different languages. We make the following contributions:
Figure 1: IndicSuperTokenizer (IST) captures superwords (e.g. “wake up”, “in the morning”) and avoids fragmenting Indic words (see for e.g. Bengali, Tamil).
  • We present IndicSuperTokenizer, a state-of-the-art tokenizer for Indic LLMs, systematically benchmarking it against popular multilingual baselines.
  • We study the impact of vocabulary size, training data, and language-specific pre-tokenization choices on fertility score, showing that careful pre-tokenization outweighs naive vocabulary scaling.
@@
  • To the best of our knowledge, we are the first to carry out a comprehensive benchmarking of a tokenizer across multiple intrinsic quality measures, as well as to study its downstream impact on task performance and LLM inference efficiency in both pretraining from scratch as well as continual pretraining settings.
@@ 2 Related Work
Tokenization strategies differ in both theory and practice. While alternate sub-word tokenization algorithms have been explored in the past such as WordPiece (song2021fast), Unigram LM (kudo2018sentencepiece), Byte Pair Encoding (BPE) remains the most widely adopted. Originally developed for compression (gage1994new) and later adapted for neural MT (sennrich2016neural), BPE merges frequent character pairs to balance coverage with efficiency. Its variants aim to address inefficiencies: PickyBPE (chizhov2024pickybpe) discards uninformative merges to improve vocabulary utility, while Scaffold-BPE (lian2024scaffold) iteratively prunes low-frequency scaffold tokens to reduce imbalance and enhance downstream performance. Recent extensions like SuperBPE (liu2025superbpe) expand beyond word boundaries, jointly learning subwords and multi-word “superwords” yielding improved compression and inference efficiency in a 2-stage curriculum. BoundlessBPE (schmidt2024boundless), another contemporary work, relaxes the Pre-tokenization word boundary constraint in an unified learning step. Our work compares these two recent approaches and builds upon SuperBPE that preserves subword coverage while capturing larger semantic units in morphologically rich Indian languages.
Multilingual Tokenizers.
Multilingual tokenization faces challenges from script diversity, morphology, and structural variation. Comparative studies show that vocabulary size and construction strategies strongly affect performance for morphologically rich languages (karthika2025multilingual), while inefficiencies in underrepresented ones, such as Ukrainian, translate to higher fertility and computational costs (maksymenko2025tokenization). Tokenization also influences how multilingual models encode morphology, as demonstrated in mT5 vs. ByT5 (dang2024tokenization). For Indic languages, tailored resources (kakwani2020indicnlp) and IndicBERT (ai4bharat2022indicbert) highlight the value of domain-specific tokenization. Recent benchmarks further reveal economic implications, with BLOOM’s tokenizer achieving the best cost efficiency among popular multilingual LLMs (adasci2024multilingual). Together, these studies show that current multilingual tokenizers fragment low-resource and morphologically rich languages, motivating approaches like ours that combine normalization, language-tailored pre-tokenization, and multi-word learning to achieve better efficiency and fairness in Indic languages. Tokenization for Indic languages presents unique challenges due to their linguistic diversity, rich morphology, and script multiplicity.
@@
Pre-tokenization plays a pivotal role in shaping token boundaries, directly influencing both compression efficiency and reasoning performance (xue2024getting). SentencePiece (kudo2018sentencepiece) introduced a language-agnostic approach by treating input as raw streams, effective for languages without whitespace boundaries. More recent approaches like BoundlessBPE (schmidt2024boundless) relax pre-token constraints to improve frequency distributions, while regex-based designs continue to prove crucial for capturing script-specific structures.
3 IndicSuperTokenizer (IST)
Language modeling involves estimating the probability distribution over text sequences, $P(S)$, where $S$ may represent a sentence, paragraph, or document. To achieve this, the text is first converted into a sequence of discrete tokens through a tokenization function $g(S)=X=(X_{1},X_{2},\ldots,X_{n})\in V^{n}$, where $V$ denotes the vocabulary and $n$ the sequence length. Tokenizers can be open-vocabulary, ensuring any string can be represented (e.g., byte-level), or closed-vocabulary, where unseen text maps to an out-of-vocabulary symbol (e.g., word lists) (rae2021scaling). In our work, we adopt an open-vocabulary approach that combines byte-pair encoding (BPE) with a UTF-8 byte fallback, following  radford2018improving. In this section, we describe our tokenizer training and evaluation approach.
3.1 Tokenizer Training
With the aim of improving fertility in Indic languages and scripts, we follow the curriculum principles as in the SuperBPE algorithm (liu2025superbpe). Specifically, we have:
Stage 2 (Superword Learning): After reaching $t$, training resumes without whitespace constraints, allowing BPE to merge across word boundaries. This enables the formation of superwords, frequent multiword expressions or collocations (e.g., “one of the”, “number of”), improving compression and reducing token counts for common phrases.
This two-stage tokenizer training is particularly effective for morphologically rich languages and scripts with complex variations where, meaningful subwords are first anchored and then composed into frequent multiword units.
3.2 Pre-tokenization
Pre-tokenization segments raw text before subword learning to improve token consistency and efficiency. We combine regex-based, Unicode normalization, and morphology-aware strategies. Unicode-aware regex separates punctuation, handles numeric groups, and aligns tokens with semantic units. NFKC normalization standardizes visually identical characters, reducing sparsity (Table 16 illustrates the effect of normalization). Morphology-aware segmentation decomposes words into roots and affixes to capture recurring morphemes. While we experimented with morphology-aware segmentation, including them in tokenization without impacting the latency is non-trivial (Refer to Appendix C.2 for details). In contrast to SuperBPE, in our Stage 1 pre-tokenization step we replace GPT-2 rules with LLaMA-4 regex for script-agnostic segmentation, improving token-to-word ratios by 38–40% (See Table LABEL:tab:regex) on Indic scripts. Stage 2 relaxes whitespace constraints to form multiword tokens capturing collocations and idioms. This design produces a script-robust tokenizer that efficiently supports multiword learning across English and Indic languages. However, unconstrained merging risks producing tokens that cross sentence boundaries,which destabilizes generation and distorts end-of-sentence probabilities. To mitigate this, we introduce sentence-level boundary constraints: merges are free within sentences but are disallowed across sentence delimiters.
Table 1: Fertility scores showing LLaMA-4 regex outperforms GPT-2 in stage-1 tokenizer training.
@@ 3.3 Training data and Vocabulary
Training a multilingual tokenizer involves careful design choices on the vocabulary size, language (or language script)-wise vocabulary distribution, and training data mix. We evaluate different vocabulary allocation strategies (Section 4.4) and conduct detailed ablations (Section 5) to inform these design choices. The final IndicSuperTokenizer that we train uses a shared vocabulary of 200200K tokens, distributed across language scripts (Figure 2), and is trained on 1010GB of multilingual high quality data curated from OLMo(olmo20252olmo2furious), Wikipedia [^2] https://en.wikipedia.org/wiki/, books, PDFs, Common Crawl, and the Sangraha dataset (khan2024indicllmsuite).
Figure 2: Vocabulary size distribution across language scripts. See Appendix A.1 for script details.
3.4 Baselines
We benchmark against 9 tokenizers, comprising: i) Indic-focused tokenizers: tokenizers designed primarily for Indian languages: Sutra (tamang2024evaluating) and Sarvam-2B (sarvam2024sarvam1) (referred as Sarvam). ii) Good Indic support tokenizers: multilingual tokenizers with demonstrated capabilities for Indic languages: Gemma-3-27B-it (team2025gemma) (referred as Gemma-3), GPT-oss (openai2025gptoss120bgptoss20bmodel) and LLaMA-4 (meta2025llama4). iii) General tokenizers: tokenizers of widely-used general-purpose LLMs: Qwen3-32B (yang2024qwen3) (referred as Qwen-3), LLaMA-3.2-1B (dubey2024llama3), Mistral-Nemo (jiang2024mistralnemo) and DeepSeek-R1 (deepseek2025deepseeekr1).
3.5 Metrics
We employ four intrinsic metrics capturing different aspects of token efficiency and informativeness: (i) Fertility score (rust2021good; scao2022bloom), measuring vocabulary granularity; (ii) Normalized Sequence Length (NSL) (dagan2024getting), quantifying sequence compression relative to a base tokenizer; (iii) Rényi’s entropy and efficiency (zouhar2023tokenization), assessing information density; and (iv) Bytes per token (kocetkov2022stack), reflecting memory and storage efficiency. We report micro-average per line at the language level. More details on the metrics and definitions in Section E in Appendix.
3.6 Evaluation Framework
We construct an evaluation set spanning 22 Indic languages, English, and code, curated from the same sources as the training corpus. Table 2 reports dataset statistics: text volume, number of lines, and average words per line per language. All metrics are computed at the line level and aggregated to the language level.
We also develop a modular evaluation framework supporting HuggingFace [^3] https://github.com/huggingface/tokenizers, SentencePiece [^4] https://github.com/google/sentencepiece, and TikToken [^5] https://github.com/openai/tiktoken tokenizers along with a comprehensive set of intrinsic metrics, including, Fertility score, normalized sequence length (NSL), Rényi entropy and efficiency, and bytes per token. We will release both the evaluation dataset and the framework for reproducible benchmarking and fair comparison of multilingual tokenizers.
Table 2: Evaluation corpus statistics across 22 Indic languages, English, and code. We report standard ISO codes here. See Section A.1 for the actual language name.
@@ 3.6 Evaluation Framework
4 Experiments and Results
4.1 Intrinsic Evaluation of Tokenizers
We achieve SOTA performance across 9 tokenizers for fertility score in consideration (see Table 3 for Indic focused or good Indic support tokenizers and an extended version Table 22 in Appendix for the rest). As shown in Table 3, IndicSuperTokenizer consistently achieves the lowest ratios across all evaluated languages, which reflects the degree of fragmentation. Bytes-per-token (Appendix  E.2) measures the average raw text bytes per token, indicating information density and sequence compactness. Table 7 shows that IndicSuperTokenizer achieves consistently higher values across languages. See Appendix  E.2 for details.
Table 3: Fertility score ($\downarrow$) comparison for Indic focused and Good support tokenizers across languages here. IST performs best in 20 of 24 languages. An extended version in Table 22 (Appendix).
@@ 4.1 Intrinsic Evaluation of Tokenizers
IST 1.85 1.74 2.04 1.47 1.45 1.12 2.17 1.77 1.23 1.21 2.19 1.58 2.30 2.28 1.63 1.62 1.65 1.39 2.59 3.72 1.45 2.12 1.88 1.44
Table 4: NSL score ($\downarrow$) comparison for Indic focused and Good support tokenizers across languages here. IST performs best in 23 of 24 languages. An extended version in Table 21 (Appendix).
IST 0.45 0.60 0.65 0.94 0.78 0.85 0.82 0.54 0.68 0.80 0.53 0.76 0.50 0.91 0.61 0.67 0.18 0.45 0.66 0.45 0.72 0.38 0.44 0.86
Normalized sequence length (Appendix  E.3) quantifies tokenized sequence length relative to a base tokenizer, indicating relative compression efficiency. Table 4 shows that IndicSuperTokenizer achieves shorter normalized sequences across languages. Rényi’s entropy quantifies the uncertainty of token distributions, while Rényi’s efficiency normalizes entropy by vocabulary size to assess utilization. Table 6 shows that IndicSuperTokenizer achieves superior efficiency across languages, reflecting effective and balanced token allocation.
Table 5: Inference latency comparison of 1B models trained with LLaMA-4 and IST tokenizers.
Model TTFT (ms) $\downarrow$ OTPT (tokens/s) $\uparrow$
LLaMA-4 19.17 $\pm$ 0.15 117.99
IST 18.98 $\pm$ 0.36 169.42
Gemma-3 GPT-oss LLaMA-4 Sarvam Sutra IST
Entropy $\downarrow$ 20.70 20.81 21.09 20.71 20.62 20.42
Efficiency $\uparrow$ 0.22 0.19 0.14 0.21 0.23 0.28
Table 7: Bytes-per-token score ($\uparrow$) comparison for Indic focused and Good support tokenizers across languages here. IST performs best in 22 of 24 languages.
@@
IST 9.12 10.15 8.18 2.84 7.86 5.44 6.29 8.95 11.01 8.59 10.30 7.80 11.33 5.67 11.11 10.39 10.07 9.40 8.70 3.60 5.40 11.32 10.70 5.55
4.2 Extrinsic Evaluation on Downstream Tasks
We also evaluated the downstream model performance by pretraining LLaMA-3.2 1B models using two tokenizers: i) IndicSuperTokenizer, our proposed tokenizer optimized for morphologically meaningful segmentation in Indic and multilingual settings, and (ii) LLaMA-4 tokenizer, chosen for comparable vocabulary size and widespread use. Both models were trained on the same dataset in iso-compute setting to ensure a fair comparison. More details in the Appendix B. We find that our tokenizer shows competitive performance across the English and Indic benchmarks.
The pretraining corpus (Table 20 in Appendix) balances coverage and domain diversity. It combines web-scale sources (Nemotron CC) for general context with structured data including MegaMath, StackV2, synthetic generations, and books. Indic-language content constitutes roughly $20\%$ of the corpus, drawn from Indic CC, Wikipedia, and Sangraha Verified, providing sufficient signal to evaluate cross-lingual and morphologically rich representation quality.
Table 8: Performance comparison of English (left) and Indic benchmarks (right).
English Benchmarks Indic Benchmarks
Dataset LLaMA-4 IST Dataset LLaMA-4 IST
HellaSwag 0.353 0.357 Indic COPA 0.544 0.556
CommonsenseQA 0.206 0.204 Indic Sentiment 0.524 0.551
OpenBookQA 0.216 0.218 Indic XNLI 0.347 0.346
Winogrande 0.504 0.510 Indic Paraphrase 0.534 0.539
GSM8K 0.016 0.018 MILU (Indic Multi-turn LU) 0.261 0.258
ARC Easy 0.623 0.630 ARC Challenge (Indic) 0.236 0.244
ARC Challenge 0.291 0.292 TriviaQA (Indic) 0.268 0.262

### v2

# MUTANT: A Recipe for Multilingual Tokenizer Design

@@ Abstract
Tokenizers play a crucial role in determining the performance, training efficiency, and the inference cost of Large Language Models (LLMs). Designing effective tokenizers for multilingual LLMs is particularly challenging due to diverse scripts and rich morphological variation. While subword methods like Byte Pair Encoding (BPE) are widely adopted, their effectiveness in multilingual settings remains underexplored. We present MUTANT, a recipe for building multilingual tokenizers, with careful vocabulary and training data design, language-aware pre-tokenization, and subword and multiword aware training. We also introduce MUTANT-Indic, a tokenizer for India-specific multilingual LLMs, that produces linguistically coherent tokens and achieves state-of-the-art performance. Evaluated across English, $22$ Indian languages and code data, our tokenizer improves the average fertility score by $39.5\%$ over LLaMA4 and by $18\%$ over Sutra (the current best). This translates to $44\%$ improvement in inference throughput over LLaMA4 while maintaining comparable performance on English and Indic benchmarks. We present detailed ablations across tokenizer training data size, vocabulary size, merging techniques, and pre-tokenization strategies, demonstrating the robustness of our design choices.
MUTANT: A Recipe for Multilingual Tokenizer Design
Souvik Rana, Arul Menezes, Ashish Kulkarni, Chandra Khatri, Shubham Agarwal Krutrim AI, Bangalore, India {souvik.rana, ashish.kulkarni, shubham.agarwal1}@olakrutrim.com
@@ 1 Introduction
Large Language Models (LLMs) (Touvron and others, 2023; Grattafiori et al., 2024; Abdin et al., 2025; Guo et al., 2025; Yang et al., 2025; Gemma Team, 2025) rely on the crucial step of tokenization, the process of converting raw text into discrete units called tokens. Among the many proposed approaches, subword tokenization schemes such as BPE (Sennrich et al., 2016a), Unigram (Kudo, 2018), WordPiece (Song et al., 2021), and their byte-level extensions have become the de facto choice. However, tokenization remains an understudied topic within the LLM literature (Dagan et al., 2024; Mielke et al., 2021), especially in multilingual settings (Petrov et al., 2023), where, skewed fertility scores across languages, often lead to concerns around fairness, high inference latency, cost and context size. For instance, with the 22 languages listed in the Eighth Schedule of the Constitution of India [^1] https://en.wikipedia.org/wiki/Languages_with_official_recognition_in_India, these issues are especially pronounced for Indic languages comprising multiple scripts and a rich morphology. A key metric for evaluating tokenizers is the “fertility score” (or token-to-word ratio) (Ali et al., 2024) where, a lower fertility score is desirable due to more efficient (and hence cheaper) LLM training and inference. Our analysis suggests that tokenizers of popular multilingual LLMs (Gemma Team, 2025; OpenAI, 2025; Meta, 2025), largely designed for English, could produce fertility scores as high as 10.5 (LLaMA-4 tokenizer for Oriya; Table 3) for Indic languages, far worse than the near-ideal scores achieved for English. This leads to longer token sequences, higher compute overheads, and poor alignment with linguistic units.
Figure 1: MUTANT: Multilingual tokenizer training recipe. Raw text is filtered and consolidated before a two-stage training process consisting of subword and multi-word learning with ablations to identify efficient design choices.
Designing an efficient tokenizer involves making careful choices around the the tokenization algorithm, size of the vocabulary (of tokens) as well as tokenizer training data. In this work, we address five core questions and investigate how to design effective multilingual tokenizers by i) examining trade-offs between low- and high-resource languages, ii) joint versus language-specific training, iii) multilingual data balancing, iv) the role of pre-tokenization, and v) the benefits of incorporating multi-word expressions via curriculum training.
@@
Supported by controlled experiments and systematic ablations, we present MUltilingual Tokenizer optimization ANd Training (MUTANT) as a practical and reproducible recipe (see Figure 1) for building equitable, efficient, and linguistically grounded multilingual tokenizers for LLMs. We apply our recipe to train MUTANT-Indic, a state-of-the-art tokenizer for Indic LLMs, demonstrating the benefits of our methodology in a real-world multilingual setting. MUTANT-Indic combines linguistically grounded pre-tokenization with a two-stage subword–multi-word learning process (Liu et al., 2025b), yielding a compact and semantically faithful vocabulary. Figure 3 illustrates some examples where our approach avoids fragmenting common words or idiomatic phrases into unnatural subunits across different languages. We make the following contributions:
  • We propose MUTANT, a principled and general recipe for training optimal multilingual tokenizers. We systematically study the effect of vocabulary size, data distribution, language-specific pre-tokenization and multi-stage training on multilingual performance.
  • We demonstrate the effectiveness by training MUTANT-Indic, that achieves state-of-the-art (SOTA) performance for Indic languages.
@@
  • To the best of our knowledge, we are the first to carry out a comprehensive benchmarking of tokenizers across multiple intrinsic and downstream LLM performance measures in both pretraining from scratch as well as continual pretraining settings. We will publicly release the evaluation framework to enable reproducibility and community benchmarking.
@@ 2 Related Work
Multilingual Tokenizers.
Multilingual tokenization faces challenges from script diversity, morphology, and structural variation. Comparative studies show that vocabulary size and construction strategies strongly affect performance for morphologically rich languages (Karthika et al., 2025a), while inefficiencies in underrepresented ones, such as Ukrainian, translate to higher fertility and computational costs (Maksymenko and Turuta, 2025). Tokenization also influences how multilingual models encode morphology, as demonstrated in mT5 vs. ByT5 (Dang et al., 2024). For Indic languages, tailored resources (Kakwani and others, 2020) and IndicBERT (AI4Bharat, 2022) highlight the value of domain-specific tokenization. Recent benchmarks further reveal economic implications, with BLOOM’s tokenizer achieving the best cost efficiency among popular multilingual LLMs (ADA Sci, 2024). Together, these studies show that current multilingual tokenizers fragment low-resource and morphologically rich languages, motivating approaches that combine normalization, language-tailored pre-tokenization, and multi-word learning to achieve better efficiency and fairness.
0250002500050000500007500075000100000100000125000125000150000150000175000175000200000200000Vocabulary Sizearbndevaengukamlorpasattate
Figure 2: Vocabulary size distribution across language scripts. See Appendix A for script details.
@@
Tokenization strategies differ in both theory and practice. While alternate sub-word tokenization algorithms have been explored in the past such as WordPiece (Song et al., 2021), Unigram LM (Kudo and Richardson, 2018), Byte Pair Encoding (BPE) remains the most widely adopted. Originally developed for compression (Gage, 1994) and later adapted for neural MT (Sennrich et al., 2016b), BPE merges frequent character pairs to balance coverage with efficiency. Its variants aim to address inefficiencies: PickyBPE (Chizhov et al., 2024) discards uninformative merges to improve vocabulary utility, while Scaffold-BPE (Lian et al., 2024) iteratively prunes low-frequency scaffold tokens to reduce imbalance and enhance downstream performance. Recent extensions like SuperBPE (Liu et al., 2025a) expand beyond word boundaries, jointly learning subwords and multi-words yielding improved compression and inference efficiency in a 2-stage curriculum. BoundlessBPE (Schmidt et al., 2024), another contemporary work, relaxes the Pre-tokenization word boundary constraint in an single stage learning step. Our work compares these two recent approaches and shows that two-stage curriculum preserves subword coverage while capturing larger semantic units in morphologically rich Indian languages.
@@
Pre-tokenization plays a pivotal role in shaping token boundaries, directly influencing both compression efficiency and reasoning performance (Xue and others, 2024). SentencePiece (Kudo and Richardson, 2018) introduced a language-agnostic approach by treating input as raw streams, effective for languages without whitespace boundaries. More recent approaches like BoundlessBPE (Schmidt et al., 2024) relax pre-token constraints to improve frequency distributions, while regex-based designs continue to prove crucial for capturing script-specific structures.
3 MUTANT: Tokenizer Training Recipe
Language modeling involves estimating the probability distribution over text sequences, $P(S)$, where $S$ may represent a sentence, paragraph, or document. To achieve this, the text is first converted into a sequence of discrete tokens through a tokenization function $g(S)=X=(X_{1},X_{2},\ldots,X_{n})\in V^{n}$, where $V$ denotes the vocabulary and $n$ the sequence length. Tokenizers can be open-vocabulary, ensuring any string can be represented (e.g., byte-level), or closed-vocabulary, where unseen text maps to an out-of-vocabulary symbol (e.g., word lists) (Rae et al., 2021). In our work, we adopt an open-vocabulary approach that combines byte-pair encoding (BPE) with a UTF-8 byte fallback, following  Radford et al. (2018). In this section, we present a systematic recipe for designing multilingual tokenizers that jointly addresses data curation, pre-tokenization, and subword–multiword learning. Figure 1 provides an end-to-end overview of the pipeline, illustrating how raw multilingual text is filtered and sampled, transformed through a fixed pre-tokenization stage, and used to train a two-stage tokenizer with explicit design choices on vocabulary size and allocation, detailed below.
@@ 3.1 Training Data and Vocabulary
The composition of the training data and the vocabulary size are design choices that directly impact the quality of tokens and the tokenizer efficiency.
Training data volume: Effective tokenizer training relies on a sufficiently large carefully curated multilingual corpus; beyond moderate scale, further data scaling yields diminishing returns (Reddy et al., 2025) in quality (Section 5).
@@
Table 1: Evaluation corpus statistics across 22 Indic languages, English, and code in Section 3.4 including size (in MB), number of lines (in K) and average words per line. We report standard ISO codes here (Appendix Section A).
Data quality and diversity: We combine diverse, high-quality sources that is representative of natural language use (Hayase et al., 2024), such as, the Web, curated multilingual datasets, Wikipedia, math, code and others. The raw data is further subjected to filtering for language/script mixing, normalization, short documents (or lines) and controlled source-wise sampling (see Appendix B.1 for MUTANT-Indic).
Vocabulary size and multilingual allocation: In multilingual settings, vocabulary allocation must be explicitly language (or script)-aware to preserve subword and multi-word granularity across both high- and low-resource languages. We determine script-wise vocabulary budgets through ablations that vary allocation across scripts, identifying distributions that balance fragmentation and coverage (Figure 2). To realize these target allocations in practice, we evaluate two strategies: i) explicit merging, where we train script-specific tokenizers and concatenates their vocabularies via rule stacking; and ii) corpus-driven alignment, where we train a single tokenizer on a multilingual corpus whose data proportions are aligned with the desired script-wise vocabulary budgets.
Explicit merging introduces distributional interference and inconsistent segmentation across scripts (Table 10). In contrast, corpus-driven alignment allows the learned vocabulary to naturally converge toward the target script-wise proportions, better mirroring corpus composition (Table 9) and achieving consistently lower fertility across scripts (Table 3), outperforming both explicit merging and public baselines.
3.2 Tokenizer Training
We adopt a two-stage curriculum following the framework of Liu et al. (2025a). This procedure forms the training strategy used in MUTANT and its Indic instantiation, MUTANT-Indic.
Table 2: Fertility scores ($\downarrow$) showing LLaMA-4 regex outperforms GPT-2 in Stage-1 tokenizer training.
@@ 3.2 Tokenizer Training
MUTANT-Indic 1.85 1.74 2.04 1.47 1.45 1.12 2.17 1.77 1.23 1.21 2.19 1.58 2.30 2.28 1.63 1.62 1.65 1.39 2.59 3.72 1.45 2.12 1.88 1.44
Table 3: Fertility score ($\downarrow$) comparison for Indic focused and good Indic support tokenizers across languages. MUTANT-Indic performs best in 20 of 24 languages. An extended version in Table 23 (Appendix).
@@
MUTANT-Indic 9.12 10.15 8.18 2.84 7.86 5.44 6.29 8.95 11.01 8.59 10.30 7.80 11.33 5.67 11.11 10.39 10.07 9.40 8.70 3.60 5.40 11.32 10.70 5.55
Table 4: Bytes-per-token score ($\uparrow$) comparison for Indic focused and good support tokenizers across languages here. MUTANT-Indic performs best in 22 of 24 languages.
Stage 2 (Multi-word Learning): After reaching $t$, training resumes without whitespace constraints, allowing BPE to merge across word boundaries. This enables the formation of multi-words, frequent multiword expressions or collocations (e.g., “in the morning” in Figure 3) improving compression and reducing token counts for common phrases.
Figure 3: MUTANT-Indic (MI) captures multi-words (e.g. “wake up", “in the morning") and avoids fragmenting Indic words (see for e.g. Bengali, Tamil).
This two-stage tokenizer training is particularly effective for scripts with complex variations where, meaningful subwords are first anchored and then composed into frequent multiword units.
3.3 Pre-tokenization
As illustrated in Figure 1, pre-tokenization is a fixed, upstream component of our tokenizer design. It segments raw text before subword learning to improve token consistency and efficiency. For MUTANT-Indic, we explore regex-based, Unicode normalization, and morphology-aware strategies. Unicode-aware regex separates punctuation, handles numeric groups, and aligns tokens with semantic units, while NFKC normalization standardizes visually identical characters to reduce orthographic sparsity (Table 25). To improve robustness across Indic scripts, we explore GPT-2 pre-tokenization rules against LLaMA-4 regex in Stage 1, where the latter yields substantially lower token fragmentation and improving token-to-word ratios by 38–40% on Indic languages (Table 2).
Script-agnostic segmentation (as in LLaMA-4 regex rules) across Indic as well as non-Indic scripts helps reduce token fragmentation, while relaxing whitespace constraints in Stage 2 enables the learning of frequent multi-word expressions (see Appendix C.7). However, unconstrained merging can destabilize generation by creating sentence-crossing tokens. Enforcing additional sentence-level boundary constraints in Stage 2 regex prevents this while preserving multi-word modeling benefits.
Additionally, morphology-aware segmentation decomposes words into roots and affixes to capture recurring morphemes. Although we explore morphology-aware segmentation, integrating it into the tokenization pipeline without incurring latency overhead is challenging (Appendix C.2).
MUTANT-Indic 0.45 0.60 0.65 0.94 0.78 0.85 0.82 0.54 0.68 0.80 0.53 0.76 0.50 0.91 0.61 0.67 0.18 0.45 0.66 0.45 0.72 0.38 0.44 0.86
Table 5: NSL score ($\downarrow$) comparison for Indic focused and Good Indic support tokenizers across languages here. MUTANT-Indic performs best in 23 of 24 languages. An extended version in Table 22 (Appendix).
3.4 Evaluation Framework
We develop a modular evaluation framework supporting HuggingFace [^2] https://github.com/huggingface/tokenizers, SentencePiece [^3] https://github.com/google/sentencepiece, and TikToken [^4] https://github.com/openai/tiktoken tokenizers along with a comprehensive set of intrinsic metrics (Appendix D), including Fertility score, Normalized Sequence Length (NSL), Rényi entropy and efficiency, and Bytes Per Token (BPT). All metrics are computed at the line level and aggregated to the language level. To assess tokenizer behavior in Indic use cases, we construct an evaluation set spanning 22 Indic languages, English, and code. Table 1 reports dataset statistics: text volume, number of lines, and average words per line per language. We will release both the evaluation dataset and the framework for reproducible benchmarking and fair comparison of multilingual tokenizers.
4 Experiments and Results
In our work, we compare MUTANT-Indic, trained using our recipe, against 9 tokenizers, comprising: i) Indic-focused tokenizers: tokenizers designed primarily for Indian languages: Sutra (Tamang and Bora, 2024) and Sarvam-2B (Sarvam, 2024) (referred as Sarvam). ii) Good Indic support tokenizers: multilingual tokenizers with demonstrated capabilities for Indic languages: Gemma-3-27B-it (Gemma Team, 2025) (referred as Gemma-3), GPT-OSS (OpenAI, 2025) and LLaMA-4 (Meta, 2025). iii) General tokenizers: tokenizers of widely-used general-purpose LLMs: Qwen3-32B (Qwen, 2024) (referred as Qwen-3), LLaMA-3.2-1B (Dubey and others, 2024), Mistral-Nemo (Mistral, 2024) and DeepSeek-R1 (DeepSeek, 2025).
4.1 Intrinsic Evaluation of Tokenizers
We evaluate tokenization quality using four complementary intrinsic metrics that capture efficiency, granularity, and information utilization: (i) Fertility score (Rust et al., 2021; Scao et al., 2022), measuring subword fragmentation; (ii) Normalized Sequence Length (NSL) (Dagan et al., 2024), quantifying relative compression against a base tokenizer; (iii) Rényi entropy and efficiency (Zouhar et al., 2023), assessing information density and vocabulary utilization; and (iv) Bytes per token (Kocetkov et al., 2022), reflecting memory and storage efficiency. All metrics are reported as micro-averages per line at the language level. Formal definitions are provided in Appendix D.
As shown in Table 3, MUTANT-Indic achieves state-of-the-art fertility scores across most languages, consistently yielding the lowest token-to-word ratios among nine tokenizers with strong Indic support (see Appendix Table 23 for extended comparisons). This indicates substantially reduced fragmentation, particularly for morphologically rich scripts. We also evaluate our Stage-1 tokenizer (trained with BPE on the same data and LLaMA-4 pretokenization). As shown in Table 24 in Appendix, our Stage-1 tokenizer, already achieves state-of-the-art (SOTA) performance across most Indic languages. Bytes-per-token results (Table 4) show that MUTANT-Indic achieves higher values across languages, reflecting more information-dense tokens and improved sequence compactness. Normalized Sequence Length scores (Table 5) further confirms that MUTANT-Indic produces shorter tokenized sequences relative to baseline tokenizers, indicating superior compression efficiency. Finally, Rényi entropy and efficiency results (Table 7) demonstrate that MUTANT-Indic achieves consistently higher efficiency across languages, reflecting effective and balanced utilization of the vocabulary. Collectively, these intrinsic results establish MUTANT-Indic as a robust and efficient tokenizer across diverse scripts and languages.
Dataset LLaMA-4 MUTANT-Indic
English Benchmarks
HellaSwag 0.353 0.357
CommonsenseQA 0.206 0.204
OpenBookQA 0.216 0.218
Winogrande 0.504 0.510
GSM8K 0.016 0.018
ARC Easy 0.623 0.630
ARC Challenge 0.291 0.292